# Task Matters: Knowledge Requirements Shape LLM Responses to Context–Memory Conflict

### v2

@@ Abstract
Large Language Models require both contextual knowledge and parametric memory, but these sources can disagree. Prior investigations on contextual question answering tasks report a preference toward parametric knowledge under conflict, yet they focus almost exclusively on tasks that should always rely on the given passage, leaving open how this behavior manifests when tasks demand different amounts and kinds of knowledge. We study this question with a model-agnostic diagnostic framework that (i) automatically detects disagreements between a model’s beliefs and a curated knowledge set, and (ii) injects controlled conflicts into tasks. The resulting datasets span two orthogonal dimensions: task knowledge reliance and conflict plausibility. Evaluating representative open-source LLMs, we find that: (1) performance degradation from conflict correlates with a task’s knowledge reliance; (2) explanatory rationales and simple reiteration both increase context reliance—helpful for context-only tasks but harmful when parametric knowledge should dominate; (3) These behaviors raise concerns about the validity of model-based evaluation and underscore the need to account for knowledge conflict in the deployment of LLMs.  [^1] Our framework and data are extensible to other models and are available at [Anonymous].
@@ 1 Introduction
Figure 1: Illustration of different evidence types that will be supplied to different tasks. In the rest of the manuscript, model internal knowledge will be referred to as No Contradiction (NC).
Figure 2: Overall diagnostic data creation flow. The lower portion is a zoom in of Evidence Creation step. After collecting the model’s parametric knowledge, the relevant support passages are further edited such that they reveal multiple levels of conflict (2. Evidence Creation) and appear in tasks that require different forms of knowledge (3. Task-Annotation).
@@
Large language models (LLMs) perform well on many knowledge-centric tasks because they encode vast amounts of parametric knowledge. In many practical settings, however, the necessary facts are supplied directly by the user in the prompt. Yet when the prompt contradicts what the model “knows,” in other words, context-memory conflict presents, LLMs frequently favor their own knowledge (Longpre et al., 2021; Chen et al., 2022; Xie et al., 2023; Jin et al., 2024a).
Prior work has quantified this bias and its impact on task performance. For instance, Longpre et al. (2021) shows that models are more likely to override the prompt when the entity is especially familiar (popular) to them, and Xie et al. (2023) finds that conflicts are resolved in favor of parametric knowledge when the contradictory context appears more plausible. These findings, however, come almost exclusively from contextual question-answering, a paradigm where the model should ground its answers strictly in the provided passage and avoid hallucinating extraneous facts. Consequently, how LLMs respond to knowledge conflict on a broader range of tasks that require different forms of knowledge utilization remains unclear Xu et al. (2024) due to the lack of data. Assessing the feasibility of a scientific idea, for instance, requires the model to (i) draw on its parametric knowledge of the broader literature, (ii) integrate novel information introduced in the prompt, and (iii) reason with both knowledge obtained in (i) and (ii). At the opposite extreme, text-copying tasks impose almost no demands on either parametric or contextual knowledge. These examples highlight that the requirement of knowledge varies sharply across tasks, underscoring the need for a systematic, task-diverse evaluation framework. We study how LLMs behave under context–memory conflict in different tasks. To this end, we introduce a systematic framework that automatically constructs diagnostic datasets across a broad spectrum of downstream tasks. Our framework identifies existing conflict between a given set of knowledge and the model being tested, and generates evaluation instances that inject controlled knowledge conflicts into different task settings. The framework spans two orthogonal dimensions: (i) task types that demand different levels of knowledge utilization, from knowledge-free (verbatim use of context only) to knowledge-intensive (reasoning with both context and prior knowledge). (ii) conflict conditions that vary in plausibility (e.g., No Contradiction NC, High/Low Plausibility Contradiction). By measuring performance differences across these conditions (fig.˜1), we can quantify the disruptive effect of knowledge conflicts on each task. The overall diagnostic data creation flow is presented in Figure 2.
Our analysis yields the following findings:
  • Through experiments on representative open-source LLMs, we find that the degree of performance degradation from knowledge conflict correlates with the task’s knowledge reliance. Conflicts barely affect tasks requiring no external knowledge, yet significantly impair knowledge-intensive tasks (§4.1).
  • Context utilization can be enhanced by using texts that contain explanatory rationales or reiteration of the same instances. However, rationales and reiteration could also bring undesired over-reliance on context when the model is expected to utilize its parametric knowledge (§4.2).
  • Our results explain findings regarding the unreliability of model-based evaluation Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025): We demonstrate that a model acting as an evaluator can be systematically biased by its own parametric knowledge, raising questions about the validity of model-based evaluation (§4.3). Conversely, we also raise the concern of susceptibility to prompt injection when the model is directed to follow the context blindly.
Finally, the framework and data we proposed can be easily extended to any current or future LLM with minimal adaptation.
@@ 2 Related Work
Xu et al. (2024) classify knowledge conflict into three categories: context-memory conflict, inter-context conflict (contradictory evidence among retrieved passages), and intra-memory conflict (inconsistent parametric beliefs), among which we focus on context-memory conflict. In this work, we focus on the context-memory conflict, which arises when a given information-bearing text chunk contradicts the model’s parametric beliefs.
@@
Methods have also been proposed to alleviate context-memory knowledge conflict. Jin et al. (2024b) identified certain attention heads that specialize in “memory” while others specialize in “context”, and therefore propose a method that dynamically prunes or patches specific attention heads that cause conflicts. Efforts have also been made to develop novel decoding methods that enhance the use of contextual knowledge Jin et al. (2024a); Shi et al. (2024); Wang et al. (2025).
@@
Most prior studies focus on contextual question answering, a setting that requires heavy reliance on the provided passages. Many other tasks, for example, grammar correction or claim verification, may need little context or, conversely, require careful integration of both parametric and contextual knowledge. This leaves the question of whether context-memory conflict poses the same impact on tasks with different knowledge demands, unanswered. To fill this gap, we keep the underlying knowledge constant while varying the task formulation, creating controlled datasets that induce different conflict levels for each target model. We introduce an analysis tool that automatically constructs model-specific test sets. Our findings indicate that both knowledge-memory conflict and blindly following the context could be particularly harmful to model-based evaluations and defending prompt injections.
@@ 3 Context-Memory Conflict Creation
Figure 2 illustrates an overview of the data construction pipeline. The process begins with identifying the pre-existing knowledge within a language model (Parametric Knowledge Collection). We use knowledge in question answering datasets that have two or more acceptable answers to one question, designed to study knowledge conflict Wan et al. (2024); Hou et al. (2024), as the knowledge source to find the stance that matches the model’s parametric belief, which will further be used to create task data. A piece of knowledge is considered part of the model’s internal belief only if the model consistently aligns with the perspective in a single answer across all prompt variations under greedy decoding, while rejecting conflicting alternatives. The variation of prompts is included in Appendix A.
With the model’s internal knowledge established, the framework generates contradictory statements based on a spectrum of conflict levels (§3.1, Evidence Creation). Leveraging these controlled contradictions, we build diagnostic datasets that consist of tasks requiring contextual knowledge, parametric knowledge, or a combination of both (§3.2, Task-Annotation). Since different models possess different parametric knowledge, the exact knowledge included in the diagnostic datasets differs by model. Each instance is then reviewed by an LLM to verify the correctness of its task type annotation (Validation).
@@ 3.1 Evidence Creation
The cognitive science literature suggests that humans address conflict between their knowledge and new information through cognitive judgment of the rationality of the concept Posner et al. (1982); Vosniadou and Brewer (1992). Xie et al. (2023) in turn suggests that LLMs could also update their answer if the provided context is convincing. We formalize it with the notion of plausibility to create a fine-grained taxonomy of conflict levels. Plausibility is defined as “at a minimum, the individual is willing to consider an alternative strategy because the recommendation is understood, coherent, and relatively simple and because the proposal is deemed a viable and logical alternative to solve the specific challenge at hand” Posner and Strike (1992). Plausibility can be used to measure how likely a human would accept new information when conflict exists. We quantify this notion by decomposing plausibility into two aspects: the content aligns with real-world or commonsense knowledge and does not violate basic logical principles. For example, suppose the model believes that grooves on the surface of Phobos, a moon of Mars, were caused by a boulder from an asteroid ejection. The conflicting statement that it was caused by gravitational pull from Mars is plausible because it conforms to common-sense knowledge. However, the idea that it was caused by a dance party is of low plausibility. With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (fig.˜1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
@@
The evidences are created following fig.˜2. Starting with an original dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2},...\},\{c_{i1},c_{i2},...\}),i\in[1,N]\}$, where $q_{i}$, $a_{i}$, $c_{i}$ corresponds to the question, answer, and context (supporting passage) of the $i$-th instance, $N$ is the size of dataset $D_{\text{orig}}$. The subscript $j$ after $i$ represents the $j$-th answer/context of the question $q_{i}$, as each question $q_{i}$ may have multiple acceptable answers. Since $D_{\text{orig}}$, coming from ConflictQA and WikiContradict, contains realistic and factually verified answers and contexts, we treat these existing answers as highly plausible. When an answer $a_{ij}$ from the original dataset contradicts the model-aligned answer $a_{ik}$ in an NC instance, we designate it as an HPC answer ($a_{i}^{\text{HPC}}=a_{ij}$), and its corresponding context as an HPC passage ($p_{i}^{\text{HPC}}=c_{ij}$). The contradicting answer $a_{ik}$ therefore becomes the NC example, namely, $a_{i}^{\text{NC}}=a_{ik}$ and $p_{i}^{\text{NC}}=c_{ik}$. To generate additional variants, we pass the passage $p_{i}^{\text{NC}}$ into an editor LLM, which is prompted to modify or rewrite it to achieve specified levels of plausibility and explanatory depth. Specifically, the editor model is instructed to rewrite the passage and degrade the plausibility while preserving contradiction to construct LPC passage $p_{i}^{\text{LPC}}$ and answer $a_{i}^{LPC}$. At the end of evidence creation, two LLMs were used to check (1) whether the passage-answer combination $(p_{i}^{\text{LPC}},a_{i}^{\text{LPC}})$ correctly answers the original question $q_{i}$; and (2) whether the generated context $p_{i}^{\text{LPC}}$ is truly low-plausibility through fact checking process.
@@ 3.2 Task Annotation
To study how models behave on tasks that require different levels of knowledge utilization, we define four tasks that differ in the extent and source of knowledge required. Examples of each task are provided in Appendix C.
@@
tasks do not require access to either contextual or parametric knowledge. We use extractive question answering as a KF task: the model is expected to extract a one-sentence answer directly from the context $p_{i}$ without engaging in reasoning, paraphrasing, or drawing upon prior knowledge. For example, the expected output in fig.˜1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context. The list of acceptable extractions is obtained and verified by GPT-4o OpenAI (2024). In the evaluation setting, the output is treated as correct as long as the extracted sentence matches one of the acceptable extractions.
@@
simulates the standard RAG setting in prior work, where models are not explicitly instructed to prioritize parametric or contextual knowledge. The model will be given two passages and is expected to answer the question based on both passages. Models are expected to acknowledge the conflict and discuss each potential answer individually. This setting naturally exposes the model to conflicts in both the context and memory.
The annotations for CK, PK, PCK, and RAG tasks derive directly from the original datasets on which our framework is built. These task types primarily differ in the number of valid answers expected and the nature of knowledge the model should rely on. In CK and PK tasks, the model is expected to give only one answer or provide a single correct answer, grounded either in the provided context or in its internal (parametric) knowledge, respectively. In PCK and RAG tasks, the model is expected to clarify that both $a_{i}^{\text{NC}}$ and the other answer are possible and explain the contradiction between the two answers.
One of the original datasets we use employs model-based evaluation to judge the correctness of free-text answers Hou et al. (2024). However, we observed that this evaluation method is susceptible to knowledge conflict, leading to inaccurate evaluations. We explore this issue further in §4.3. Therefore, we modify the non-extractive tasks to be multiple-choice questions. Each instance presents four answer options; the model must first generate an explanation, then select the most appropriate answer. To assess the performance of the target model, we report the accuracy for CK and PK tasks, F1 for KF, PCK, and RAG. To obtain high-quality texts, we use GPT-4o as the base model to create evidence and validate the diagnostic data. Then, we analyze the instruction-tuned version of Mistral-7B Jiang et al. (2023), OLMo2-7B, OLMo2-13B OLMo et al. (2024), Qwen2.5-7B, and Qwen2.5-14B Qwen et al. (2025), all of which are widely used open-weight models that represent diverse training paradigms. The resulting diagnostic data is composed of 2,893 instances for Mistral-7B, 177 instances for OLMo, and 6,217 instances for Qwen2.5-7B. Each instance includes three different evidence types (NC, HPC, LPC); thus, the resulting task data has three times the number of instances.

### v3

@@ Abstract
Large language models (LLMs) draw on both contextual information and parametric memory, yet these sources can conflict. Prior studies have largely examined this issue in contextual question answering, implicitly assuming that tasks should rely on the provided context, leaving unclear how LLMs behave when tasks require different types and degrees of knowledge utilization. We address this gap with a model-agnostic diagnostic framework that holds underlying knowledge constant while introducing controlled conflicts across tasks with varying knowledge demands. Experiments on representative open-source LLMs show that performance degradation under conflict is driven by both task-specific knowledge reliance and conflict plausibility; that strategies such as rationales or context reiteration increase context reliance, helping context-only tasks but harming those requiring parametric knowledge; and that these effects bias model-based evaluation, calling into question the reliability of LLMs as judges. Overall, our findings reveal that context–memory conflict is inherently task-dependent and motivate task-aware approaches to balancing context and memory in LLM deployment and evaluation. [^1] Our framework and data are available at [Anonymous].
@@ 1 Introduction
Figure 1: Overview of the types of contexts and tasks in our evaluation. Context types vary in the level of conflict, while the tasks impose different knowledge constraints.
@@
Large language models (LLMs) perform well on many knowledge-centric tasks because they encode vast amounts of parametric knowledge within their parameters. In many practical settings, however, the necessary facts are supplied directly by the user in the prompt, which often includes information that is updated after the model’s knowledge cutoff. Context–memory conflict arises when such input contradicts what the model “knows,” and LLMs often favor their own parametric knowledge over the provided context (Longpre et al., 2021; Chen et al., 2022; Xie et al., 2023; Jin et al., 2024a; Liu et al., 2025).
Figure 2: Overall diagnostic data creation flow. The lower portion is a zoom in of Evidence Creation step. After collecting the test model’s parametric knowledge, the supporting passages are further edited to reveal multiple levels of conflict (2. Evidence Creation) and appear in different tasks (3. Task-Annotation).
Prior work has shown that context–memory conflict substantially affects model behavior, but the resulting conclusions remain fragmented and do not yet form a coherent explanation. Some studies find that models systematically favor their parametric knowledge, particularly when conflicting entities are familiar (Longpre et al., 2021), while others report that models instead follow contextual evidence when it is sufficiently coherent or convincing (Xie et al., 2023; Jin et al., 2024a). Most of these results, however, are derived from contextual question answering, where models are expected to rely exclusively on the provided passage. As a result, it remains unclear how LLMs should behave under context–memory conflict when tasks require different forms of knowledge utilization (Xu et al., 2024).
In practice, tasks vary sharply in their reliance on contextual versus parametric knowledge. Extractive tasks, such as text copying, impose little need for prior knowledge, whereas tasks such as literature review require integrating background knowledge with new information published after the model’s knowledge cutoff. Between these extremes lie tasks that demand paraphrasing, selective grounding, or explicit reconciliation of conflicting evidence. Treating all such settings as interchangeable obscures how context–memory conflict functions across real-world applications.
In this work, we show that the impact of context–memory conflict is task-dependent. We demonstrate that performance degradation under conflict correlates with a task’s knowledge requirements, in addition to the level of conflict alone. To make this dependence explicit, we introduce a diagnostic framework that holds underlying knowledge constant while systematically varying task formulations. This allows us to isolate how the same conflict affects models when a task requires (i) no knowledge beyond surface extraction, (ii) grounding in context, (iii) reliance on parametric knowledge, or (iv) integration of both. Our framework automatically identifies a model’s parametric beliefs and injects controlled contradictions into downstream tasks (Figure˜1), producing model-specific diagnostic datasets that vary both conflict level and task knowledge requirements. Evaluating models on these datasets enables us to disentangle the effects of knowledge conflict from task demands. This allows us to quantify how knowledge conflict disrupts model behavior in a task-aware manner (Figure˜2).
Across representative LLMs, we find that the effects of context–memory conflict are strongly task-dependent: conflicts have minimal impact on tasks requiring little or no knowledge utilization, yet significantly impair knowledge-intensive tasks, even under explicit instructions. The factor of tasks explains the prior findings that would appear inconsistent if examined without the lens of task. Beyond overall performance, we find that simple strategies such as reiteration encourage context reliance, without the need to alter the inference mechanism. Reiteration improves performance when the task is intended to rely primarily on contextual information (Wang et al., 2022; Wu et al., 2024), but degrades performance when successful task completion requires dominant use of parametric knowledge. Finally, we demonstrate that studying context–memory conflict particularly matters for the task of model-based evaluation. Using LLMs as judges (Zheng et al., 2023; Liu et al., 2023; Ru et al., 2024; Chen et al., 2025) constitutes a task in which the model is required to balance parametric knowledge and contextual information, while the model’s own parametric knowledge can implicitly bias evaluations, particularly when factual correctness conflicts with the provided context. Together, our findings suggest that context–memory conflict cannot be meaningfully studied or mitigated without explicit consideration of task knowledge requirements, and motivate task-aware approaches to balancing contextual and parametric knowledge.
@@ 2 Related Work
Xu et al. (2024) classify knowledge conflict into three categories: context-memory conflict, inter-context conflict (contradictory evidence among retrieved passages), and intra-memory conflict (inconsistent parametric beliefs). We focus on the context-memory conflict, which arises when a given information-bearing text chunk contradicts the model’s parametric beliefs.
@@
Methods have also been proposed to alleviate context-memory knowledge conflict. Jin et al. (2024b) identified certain attention heads that specialize in “memory” while others specialize in “context”, and therefore propose a method that dynamically prunes or patches specific attention heads that cause conflicts. Li et al. (2025) instead propose a two-step inference intervention that has each step focus on either the context or the memory heads. Efforts have also been made to develop novel decoding methods that enhance the use of contextual knowledge Jin et al. (2024a); Shi et al. (2024); Wang et al. (2025a). More recently, work has explored post-hoc, inference-time controls that steer context reliance without fine-tuning, including proxy-model steering (Wang et al., 2025b) and test-time attention interventions (Li et al., 2025).
@@
Most prior studies focus on contextual question answering, a setting that requires heavy reliance on the provided passages. Many other tasks, for example, grammar correction or claim verification, may need little context or, conversely, require careful integration of both parametric and contextual knowledge. This leaves the question of whether context-memory conflict poses the same impact on tasks with different knowledge demands unanswered. To fill this gap, we keep the underlying knowledge constant while varying the task formulation, creating controlled datasets that induce different conflict levels for each target model. We introduce an analysis tool that automatically constructs model‑specific test sets. Our findings indicate that both knowledge-memory conflict and blindly following the context could be particularly harmful to model-based evaluations, further motivating task-dependent methods for dynamically balancing context and memory (Wang et al., 2025b; Li et al., 2025).
@@ 3 Context-Memory Conflict Creation
Figure 2 illustrates an overview of the data construction pipeline. The process begins with identifying the pre-existing knowledge within a language model (Parametric Knowledge Collection). We use the knowledge from knowledge conflict question answering datasets Wan et al. (2024); Hou et al. (2024) that have two or more acceptable answers to one question, using them to identify the stance aligned with the model’s parametric belief, which then serves as the basis for constructing task data. A piece of knowledge is considered part of the model’s internal belief only if the model consistently aligns with the perspective in a single answer across all prompt variations under greedy decoding, while rejecting conflicting alternatives, with details included in Appendix A. With the model’s internal knowledge, the framework generates contradictory statements based on a spectrum of conflict levels (§3.1, Evidence Creation). Leveraging these controlled contradictions, we build diagnostic datasets that consist of tasks requiring contextual knowledge, parametric knowledge, or a combination of both (§3.2, Task-Annotation). Two LLMs then review each instance to verify the correctness of its task type annotation, with a subsample of the instances verified by human annotators (Validation, Appendix B).
@@ 3.1 Evidence Creation
The cognitive science literature suggests that humans resolve conflicts between prior knowledge and new information by judging the rationality of alternative concepts Posner et al. (1982); Vosniadou and Brewer (1992). Similarly, Xie et al. (2023) shows that LLMs may also revise their answers when the context is sufficiently convincing. We formalize this with the notion of plausibility, defined as the willingness to consider an alternative strategy when it is understood, coherent, relatively simple, and deemed a viable solution Posner and Strike (1992). Plausibility can be used to measure how likely a human is to accept new information in the presence of conflict. We decompose it into two criteria: alignment with real-world or commonsense knowledge and consistency with basic logical principles. For example, if a model believes that grooves on Phobos were caused by a boulder from an asteroid ejection, the claim that they resulted from Mars’s gravitational pull is plausible, as it fits common-sense knowledge. By contrast, attributing the grooves to a dance party is implausible. With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (Figure˜1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
@@
The evidences are created following Figure˜2. Starting with an original dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2},...\},\{c_{i1},c_{i2},...\}),i\in[1,N]\}$, where $q_{i}$, $a_{i}$, $c_{i}$ corresponds to the question, answer, and context (supporting passage) of the $i$-th instance, $N$ is the size of dataset $D_{\text{orig}}$. The subscript $j$ after $i$ represents the $j$-th answer/context of the question $q_{i}$, as each question $q_{i}$ may have multiple acceptable answers. Since $D_{\text{orig}}$, coming from ConflictQA and WikiContradict, contains realistic and factually verified answers and contexts, we treat these existing answers as highly plausible. When an answer $a_{ij}$ from the original dataset contradicts the model-aligned answer $a_{ik}$ in an NC instance, we designate it as an HPC answer ($a_{i}^{\text{HPC}}=a_{ij}$), and its corresponding context as an HPC passage ($p_{i}^{\text{HPC}}=c_{ij}$). The contradicting answer $a_{ik}$ therefore becomes the NC example, namely, $a_{i}^{\text{NC}}=a_{ik}$ and $p_{i}^{\text{NC}}=c_{ik}$. To generate additional variants, we pass the passage $p_{i}^{\text{NC}}$ into an editor LLM, which is prompted to modify or rewrite it to achieve specified levels of plausibility and explanatory depth. Specifically, the editor model is instructed to rewrite the passage and degrade the plausibility while preserving contradiction to construct LPC passage $p_{i}^{\text{LPC}}$ and answer $a_{i}^{LPC}$. At the end of evidence creation, two LLMs were used to check (1) whether the passage-answer combination $(p_{i}^{\text{LPC}},a_{i}^{\text{LPC}})$ correctly answers the original question $q_{i}$; and (2) whether the generated context $p_{i}^{\text{LPC}}$ is truly low-plausibility through fact checking process.
@@ 3.2 Task Annotation
To study how models behave on tasks that require different levels of knowledge utilization, we define four tasks that differ in the extent and source of knowledge required. Examples of each task are provided in Appendix D.
@@
tasks do not require access to either contextual or parametric knowledge. We use extractive question answering as a KF task: the model is expected to extract a one-sentence answer directly from the context $p_{i}$ without engaging in reasoning, paraphrasing, or drawing upon prior knowledge. For example, the expected output in Figure˜1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context. The list of acceptable extractions is obtained and verified by GPT-4o OpenAI (2024). In the evaluation setting, the output is treated as correct as long as the extracted sentence matches one of the acceptable extractions.
@@
simulates the standard RAG setting, where models are not explicitly instructed to prioritize parametric or contextual knowledge. The model will be given two passages and is expected to answer the question based on both passages. Models are expected to acknowledge the conflict and discuss each potential answer individually. This setting serves as both a complementary setting to PCK tasks and creates a test bed that is closer to real-world applications, as Hagström et al. (2024) show that insights based on synthesized data are not guaranteed to generalize to real-world scenarios.
The annotations for all five tasks are derived directly from the original datasets on which our framework is built. These task types primarily differ in the number of valid answers expected and the nature of knowledge the model should rely on. In KF tasks, the model is only expected to perform extractions. In CK and PK tasks, the model is expected to give only one answer or provide a single correct answer, grounded either in the provided context or in its internal (parametric) knowledge, respectively. In PCK and RAG tasks, the model is expected to clarify that both $a_{i}^{\text{NC}}$ and the other answer are possible and explain the contradiction between the two answers.
One of the original datasets we use employs model-based evaluation to judge the correctness of free-text answers Hou et al. (2024). However, we observed that this evaluation method is susceptible to knowledge conflict, leading to inaccurate evaluations. We explore this issue further in §4.3. Therefore, we modify the non-extractive tasks to be multiple-choice questions. Each instance presents four answer options; the model must first generate an explanation, then select the most appropriate answer. To assess the performance of the target model, we report the accuracy for CK and PK tasks, F1 for KF, PCK, and RAG. To obtain high-quality texts, we use GPT-4o as the base model to create evidence and validate the diagnostic data. Then, we analyze the instruction-tuned version of Mistral-7B Jiang et al. (2023), OLMo2-7B, OLMo2-13B OLMo et al. (2024), Qwen2.5-7B, and Qwen2.5-14B Qwen et al. (2025), all of which are widely used open-weight models that represent diverse training paradigms. The resulting dataset statistics are presented in Appendix C.